OpenStudy (anonymous):

what is the quotient 6-x/x^2+2x-3 divided by x^2-4x-12/x^2+4x+3

OpenStudy (anonymous):

@mathstudent55 @JoannaBlackwelder

OpenStudy (tkhunny):

Have you considered factoring things?

OpenStudy (anonymous):

@tkhunny I don't know what to do at all. Can we walk through it? And should I post the answer choices?

OpenStudy (jenniferjuice):

yeah post the answer choices it always narrows it down

OpenStudy (tkhunny):

You walk through it. Factor these: 6-x x^2+2x-3 x^2-4x-12 x^2+4x+3

OpenStudy (anonymous):

How do I factor them?

OpenStudy (tkhunny):

What class are you in?

OpenStudy (anonymous):

x^2+4x+3 factored would be (x-1) (x-3) right?

OpenStudy (anonymous):

6-x (x-6) x^2+2x-3 (x+3) (x-1) x^2-4x-12 (x+2 (x-6) x^2+4x+3 (x-1) (x-3) @tkhunny

OpenStudy (jenniferjuice):

tkhunny just comes and goes..0.0

OpenStudy (anonymous):

Am I even right?

OpenStudy (tkhunny):

I'm confused. You say this: "How do I factor them?" Then you pop out with this: "6-x (x-6) x^2+2x-3 (x+3) (x-1) x^2-4x-12 (x+2 (x-6) x^2+4x+3 (x-1) (x-3) " So, you DID know how to do it? Why were you declining? Perfect. Now, rewrite that horrible original fraction in terms of these factored versions.

OpenStudy (anonymous):

(x-6)/(x+3)(x-1) divided by (x+2) (x-6)/(x-1)(x-3) @tkhunny

OpenStudy (tkhunny):

Nice. Now, utilize the idea of division of fractions (reciprocal and multiply) and rewrite it again.

OpenStudy (anonymous):

Wait. What does that mean? I've completely lost everything. @tkhunny

OpenStudy (tkhunny):

Calculate the quotient: \(\dfrac{\dfrac{1}{3}}{\dfrac{9}{5}}\).

OpenStudy (anonymous):

5/27 @tkhunny

OpenStudy (tkhunny):

How did you determine that?

OpenStudy (anonymous):

I used a fractions calculator

OpenStudy (tkhunny):

{sigh} That doesn't help. You should know how to divide fractions. \(\dfrac{\dfrac{1}{3}}{\dfrac{9}{5}} = \dfrac{1}{3}\cdot\dfrac{5}{9} = \dfrac{5}{27}\) Have you seen the process usually called "reciprocal and multiply"? That is all we are doing.

OpenStudy (anonymous):

I got 5/27?... And no. This unit was one of the hardest units I've ever done.

OpenStudy (tkhunny):

You don't seem to be answering my questions. You have not yet responded to the concept of "reciprocal and multiply". HAVE you seen it or heard of it? It is the way to divide fractions. (2/5) / (1/7) = (2/5)*(7/1) That has to look familiar.

OpenStudy (anonymous):

I obviously responded to your question when I said "And no". It looks semi familiar but I don't remember working it like that at all.

OpenStudy (tkhunny):

It's time for you to do it with the giant expression. (x-6)/[(x+3)(x-1)] divided by [(x+2) (x-6)]/[(x-1)(x-3)]

OpenStudy (bradely):

x^2+4x+3= (x+1) (x+3)

OpenStudy (bradely):

Please check this factorization

OpenStudy (anonymous):

Sorry. I got busy. The answer I got was 1/(x-3)(x-1)^2(x+2)(x+3)

OpenStudy (bradely):

its wrong option 3 is correct

OpenStudy (anonymous):

So its C. Can you help me with another? @bradely

OpenStudy (anonymous):

simplify the complex fraction x/x+4/1/x+1/x+4

OpenStudy (anonymous):

Now how do I do this one? :P

OpenStudy (bradely):

i will respond u

OpenStudy (anonymous):

Okay. Thanks

OpenStudy (bradely):

Option 4 is correct please check in site

OpenStudy (anonymous):

I saw it! Thanks man. You're rad B)